OpenStudy (anonymous):

4( x – 2 ) = 5( x + 3)

OpenStudy (anonymous):

distribute the 4 and 5 to each side respectively 4x - 8 = 5x + 15 combine like terms 4x - 5x = 8 + 16 solve -1x = 24 x = -24

OpenStudy (anonymous):

ah pellet, i mean 15, not 16 lmao. -1x = 23 x = -23